Circular No. 7968

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2002eq, 2002ez, AND 2002fa M. Hamuy, Carnegie Observatories, reports that a spectrum (range

380–730 nm) of SN 2002eq (cf. IAUC 7959), obtained on Sept. 12.17 UT with the Las Campanas Dupont 2.5-m telescope (+ Modular Spectrograph), shows it to be a type-Ia supernova. The spectrum bears strong resemblance to SN 1994D at 20 days past maximum; a redshift of ≈ 0.088 is derived by comparison to SN 1994D. A spectrum of SN 2002ez (cf. IAUC 7967), obtained on Sept. 12.27, bears resemblance to the type-Ia SN 1994D at 18 days past maximum; a redshift of ≈ 0.043 is derived by comparison to SN 1994D. A spectrum of SN 2002fa (cf. IAUC 7967), obtained on Sept. 12.08, shows it to be of type II with a broad Hα line exhibiting a P-Cyg profile. The minimum of the absorption component yields an expansion velocity of 8100 km/s. A spectrum of the nucleus of the host galaxy yields a redshift of 0.060.

RADIO TRANSIENT IN SCORPIUS M. P. Rupen, V. Dhawan, and A. J. Mioduszewski, National Radio

Astronomy Observatory, report that observations with the Very Large Array (VLA) on Sept. 9 revealed a new ∼ 3-mJy source at 8.46 GHz at a nominal position α = 16^{h}28^{m}47^{s}.3, δ = −41^{o}52′35′′ (equinox 2000.0). Although the detection is highly significant (11σ), the low declination, poor weather, and recent re-positioning of several of the VLA’s constituent antennas combine to make the flux density uncertain to perhaps 50 percent and the position accurate to only a few arcsec. The beam size was 3′′.6 × 0′′.9 FWHM at a position angle of −7^{o}. VLA observations of the same field over the past year, primarily at 8.46 GHz, have given numerous nondetections, to rms noise limits of 0.1–0.3 mJy/beam. Based on the rarity of strong, variable radio sources at this frequency, and the close (3′′.4) positional agreement, Rupen et al. suggest this is a radio counterpart to GSC 7861.1088 = 1RXS J162848.1–415241, proposed by Tsarevsky

et al. (cf. http://vsnet.kusastro.kyoto-u.ac.jp/vsnet/Mail/vsnet-

campaign/msg01058.html) as a probable microquasar. Further radio, optical, infrared, and x-ray observations are urgently requested.

COMET P/2002 Q1 (VAN NESS) This comet is of short period, as shown by the orbital elements on

MPEC 2002-R52 [T = 2002 July 15.99 TT, q = 1.5214 AU, e = 0.5695, ω = 184^{o}.76, Ω = 174^{o}.11, i = 36^{o}.37 (equinox 2000.0), P = 6.64 yr].

2002 September 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green